# Exact and efficient basis pursuit denoising via differential inclusions and a selection principle

### v1

@@ Appendix A Mathematical background
$\partial\chi_{C}(\boldsymbol{x})\coloneqq N_{C}(\boldsymbol{x})=\left\{\boldsymbol{s}\in\mathbb{R}^{n}:\langle\boldsymbol{s},\boldsymbol{x}-\boldsymbol{y}\rangle\geqslant 0\ \text{for every $\boldsymbol{y}\in C$}\right\}.$
The set $N_{C}(\boldsymbol{x})$ is called the normal cone of $C$ at $\boldsymbol{x}\in C$, and it is a closed convex cone.
@@
See [55, Theorem 10.1, page 422] for the proof. ∎
@@
$\partial\chi_{C}(\boldsymbol{x})=\mathrm{cone}\{\boldsymbol{u}_{j}:j\in\mathcal{E}(\boldsymbol{x})\}=\left\{\sum_{j\in\mathcal{E}(\boldsymbol{x})}\eta_{j}\boldsymbol{u}_{j}:\eta_{j}\geqslant 0\right\}.$
@@
This follows from Definitions A.2 and A.18. See [38, Example 5.2.6] for details. ∎
@@
See [39, Corollary 3.1.2] for the proof of (i), [39, Theorem 2.3.3] for the proof of (ii), and [39, Theorem 2.2.1, Theorem 2.2.3, and Theorem 3.2.1.] for the proof of (iii). The proof of (iv) follows by using (i) with $f_{1}=f$ and $f_{2}=g\circ\bm{A}$ and then using (iii). ∎
@@
    $\sup_{\boldsymbol{p}\in\mathbb{R}^{m}}\left\{-f_{1}^{*}(-\bm{A}^{*}\boldsymbol{p})-f_{2}^{*}(\boldsymbol{p})\right\}$ (60)
@@
See [54, Theorem 31.2 and Corollary 31.2.1] and [3, Proposition 1, Theorem 1 and Theorem 2, pages 163–167] for a proof. The equivalences in (59) and (61) follow from (58) and the rules in Proposition A.2. The equivalence in (62) follows from Fenchel’s inequality (56). ∎
@@
The assumptions in (58) read: there exists $\boldsymbol{x}\in\mathrm{dom~{}}f_{1}$ such that $\bm{A}\boldsymbol{x}\in\mathrm{ri}\ \mathrm{dom~{}}f_{2}$ and there exists $\boldsymbol{p}\in\mathrm{dom~{}}f_{2}^{*}$ such that $-\bm{A}^{\top}\boldsymbol{p}\in\mathrm{dom~{}}f_{1}^{*}$. If the first assumption holds, then Proposition A.2 implies the function $\boldsymbol{x}\mapsto f_{1}(\boldsymbol{x})+f_{2}(\bm{A}\boldsymbol{x})$ is in $\Gamma_{0}(\mathbb{R}^{n})$. Moreover,
@@
Appendix B Detailed Proofs
@@
Part 1. First, we show there is $\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})>0$ such that $\left\|{-\bm{A}^{\top}(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}))}\right\|_{\infty}\leqslant 1$ for every $\Delta\in[0,\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})]$. Multiply the vector $-\bm{A}^{\top}(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}))$ by the matrix of signs $\bm{D}(\boldsymbol{p}_{0})$ and take the inner product with respect to the unit vector $\boldsymbol{e}_{j}$ with $j\in\{1,\dots,n\}$:
@@
First, suppose $j\in\mathcal{E}(\boldsymbol{p}_{0})$. Then $\langle-\bm{D}(\boldsymbol{p}_{0})\bm{A}^{\top}\boldsymbol{p}_{0},\boldsymbol{e}_{j}\rangle=1$ and $\langle\bm{D}(\boldsymbol{p}_{0})\bm{A}^{\top}\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}),\boldsymbol{e}_{j}\rangle\geqslant 0$ by the KKT condition (17a). If the latter is strictly positive, then we can increase $\Delta$ until the left-hand-side of (63) is equal to $-1$. The smallest such number is
@@
[…]
@@
[…]
@@
Part 3. Let $\Delta\in[0,\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})]$ and $\boldsymbol{q}\in\mathbb{R}^{m}$. Since $-\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})\in\partial_{\boldsymbol{p}}V(\boldsymbol{p}_{0};t,\boldsymbol{b})$ and $V(\cdot;t,\boldsymbol{b})$ is $t$-strongly convex, the subdifferentiability property implies
@@
$-(1-t\Delta)\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})\in\partial_{\boldsymbol{p}}V(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b});t,\boldsymbol{b})\ \text{for every $\Delta\in[0,\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})]$}.$
@@
$-\mathrm{proj}_{\partial_{\boldsymbol{p}}V(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b});t,\boldsymbol{b})}(\boldsymbol{0})\equiv\boldsymbol{d}(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b});t,\boldsymbol{b})=(1-t\Delta)\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}).$
@@
As a consequence, we can extend the analysis done in Theorem 4.1 to the value $\tau=\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})$. That is, $\boldsymbol{p}(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}))=\boldsymbol{p}_{a}(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}))$, and so (23) holds at $\tau=\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})$.
Next, we prove inequality (28). It suffices to show
$\boldsymbol{d}(\boldsymbol{p}(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}));t,\boldsymbol{b})\neq e^{-t\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})}\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}).$
Indeed, the projection onto the closed and convex set $\partial_{\boldsymbol{p}}V(\boldsymbol{p}(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}));t,\boldsymbol{b})$ is unique (see A.17), so the subgradients are different if and only if their norms are different. First, we apply Theorem 4.1 with the initial value $\boldsymbol{p}(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}))$ at $\tau=\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})$ to find
in the time interval $[\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}),\tau_{*}(\boldsymbol{p}(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}));t,\boldsymbol{b})]$. Assume for a contradiction that $\boldsymbol{d}(\boldsymbol{p}(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}));t,\boldsymbol{b})=e^{-t\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})}\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})$. Then (68) can be developed further into
However, this would imply
$\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})+f(\tau-\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}),t)e^{-t\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})}>\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})$
for every $\tau\in(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}),\tau_{*}(\boldsymbol{p}(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}));t,\boldsymbol{b})]$, contradicting the definition of the number $\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})$ in (18). We conclude $\boldsymbol{d}(\boldsymbol{p}(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}));t,\boldsymbol{b})\neq e^{-t\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})}\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})$.
B.3 Proof of Lemma 5.1
Part 1. Use Proposition 3.1 with hyperparameter $t_{0}+\delta_{0}$ and data $\boldsymbol{b}+\boldsymbol{\delta}\boldsymbol{b}$ to get
$\boldsymbol{d}(\boldsymbol{p}_{0};t_{0}+\delta_{0},\boldsymbol{b}+\boldsymbol{\delta}\boldsymbol{b})=\bm{A}\bm{D}(\boldsymbol{p}_{0})\hat{\boldsymbol{u}}(\boldsymbol{p}_{0};t_{0}+\delta_{0},\boldsymbol{b}+\boldsymbol{\delta}\boldsymbol{b})-(\boldsymbol{b}+\boldsymbol{\delta}\boldsymbol{b})-(t_{0}+\delta_{0})\boldsymbol{p}_{0},$
@@
$\hat{\boldsymbol{u}}(\boldsymbol{p}_{0};t_{0}+\delta_{0},\boldsymbol{b}+\boldsymbol{\delta}\boldsymbol{b})\in\operatorname*{arg\,min}_{\boldsymbol{u}\in\mathbb{R}^{n}}\left\|{\bm{A}\bm{D}(\boldsymbol{p}_{0})\boldsymbol{u}-\boldsymbol{b}-\boldsymbol{\delta}\boldsymbol{b}-(t_{0}+\delta_{0})\boldsymbol{p}_{0}}\right\|_{2}^{2}$
subject $\qquad\qquad\text{subject to}\ \begin{cases}\boldsymbol{u}_{\mathcal{E}(\boldsymbol{p}_{0})}\geqslant\boldsymbol{0}\\ \boldsymbol{u}_{\mathcal{E}^{\mathsf{C}}(\boldsymbol{p}_{0})}=\boldsymbol{0}.\end{cases}$
Now let $\boldsymbol{v}\in\mathbb{R}^{m}$ and use the change of variables $\boldsymbol{u}=\hat{\boldsymbol{u}}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b})+\boldsymbol{v}$ in the optimization problem above. The constraints of the problem become $\boldsymbol{v}_{\mathcal{E}(\boldsymbol{p}_{0})}\geqslant-\hat{\boldsymbol{u}}_{\mathcal{E}(\boldsymbol{p}_{0})}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b})$ and $\boldsymbol{v}_{\mathcal{E}^{\mathsf{C}}(\boldsymbol{p}_{0})}=\boldsymbol{0}$, while the objective function becomes
$\left\|{\bm{A}\bm{D}(\boldsymbol{p}_{0})\boldsymbol{u}-\boldsymbol{b}-\boldsymbol{\delta}\boldsymbol{b}-(t_{0}+\delta_{0})\boldsymbol{p}_{0}}\right\|_{2}^{2}$
$\qquad\qquad\qquad=\left\|{\bm{A}\bm{D}(\boldsymbol{p}_{0})\hat{\boldsymbol{u}}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b})+\bm{A}\bm{D}(\boldsymbol{p}_{0})\boldsymbol{v}-\boldsymbol{b}-\boldsymbol{\delta}\boldsymbol{b}-(t_{0}+\delta_{0})\boldsymbol{p}_{0}}\right\|_{2}^{2}$
$\qquad\qquad\qquad=\left\|{\left(\bm{A}\bm{D}(\boldsymbol{p}_{0})\hat{\boldsymbol{u}}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b})-\boldsymbol{b}-t\boldsymbol{p}_{0}\right)+\bm{A}\bm{D}(\boldsymbol{p}_{0})\boldsymbol{v}-\boldsymbol{\delta}\boldsymbol{b}-\delta_{0}\boldsymbol{p}_{0}}\right\|_{2}^{2}$
$\qquad\qquad\qquad=\left\|{\boldsymbol{d}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b})+\bm{A}\bm{D}(\boldsymbol{p}_{0})\boldsymbol{v}-\boldsymbol{\delta}\boldsymbol{b}-\delta_{0}\boldsymbol{p}_{0}}\right\|_{2}^{2},$
From this, we obtain (31) and (32).
Part 2. Set $\boldsymbol{d}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b})=\boldsymbol{0}$, $\delta_{0}=t-t_{0}$ with $t\in[0,t_{0}]$ and $\boldsymbol{\delta}\boldsymbol{b}=-(t_{0}-t)\boldsymbol{q}$ in Part 1, then factor out the term $(1-t/t_{0})$ outside the optimization problem to obtain immediately (33) and (34). Note that the case $t_{0}=0$ works because we assume $t\in[0,t_{0}]$.
B.4 Proof of Proposition 5.1
Part 1. First, we invoke Lemma 5.1(ii) with
@@
$\boldsymbol{d}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})$ $;t,\boldsymbol{b}-(t_{0}-t)\boldsymbol{q})=$ (69)
$(1-t/t_{0})(\bm{A}\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\hat{\boldsymbol{v}}(t_{0},t,\boldsymbol{b},\boldsymbol{q})+t_{0}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})+\boldsymbol{q})),$
@@
$\hat{\boldsymbol{v}}(t_{0},t,\boldsymbol{b},\boldsymbol{q})\in\operatorname*{arg\,min}_{\boldsymbol{v}\in\mathbb{R}^{n}}\left\|{\bm{A}\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\boldsymbol{v}+t_{0}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})+\boldsymbol{q})}\right\|_{2}^{2}$ (70)
subject $\text{subject to}\ \begin{cases}&\boldsymbol{v}_{j}\geqslant-|\boldsymbol{x}^{s}_{j}(t_{0},\boldsymbol{b})|/(1-t/t_{0})\ \text{if $j\in\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))$ and $\boldsymbol{x}^{s}_{j}(\boldsymbol{b},t_{0})\neq 0$}\\ &\boldsymbol{v}_{j}\geqslant 0\ \text{if $j\in\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))$ and $\boldsymbol{x}^{s}_{j}(\boldsymbol{b},t_{0})=0$,}\\ &\boldsymbol{v}_{\mathcal{E}^{\mathsf{C}}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}=\boldsymbol{0}.\end{cases}$
At $t=t_{0}$, problem (70) reduces to (35) with the identification $\hat{\boldsymbol{v}}(t_{0},t_{0},\boldsymbol{b},\boldsymbol{q})\equiv\hat{\boldsymbol{v}}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})$.
We will now identify how the solution $\hat{\boldsymbol{v}}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})$ behaves as we decrease $t_{0}$. There are two potential sources of changes: the set of constraints in (70) and the minimal selection principle via the evolution rule (20) in Proposition 3.1. We turn to these two sources in turn.

### v2

@@ Appendix A Mathematical background
$\partial\chi_{C}(\boldsymbol{x})\coloneqq\left\{\boldsymbol{s}\in\mathbb{R}^{n}:\langle\boldsymbol{s},\boldsymbol{x}-\boldsymbol{y}\rangle\geqslant 0\ \text{for every $\boldsymbol{y}\in C$}\right\}.$
This set is called the normal cone of $C$ at $\boldsymbol{x}\in C$, and it is a closed convex cone.
@@
See [43, Theorem 10.1, page 422] for the proof. ∎
@@
$\partial\chi_{C}(\boldsymbol{x})=\left\{\sum_{j\in\mathcal{E}(\boldsymbol{x})}\eta_{j}\boldsymbol{u}_{j}:\eta_{j}\geqslant 0\right\}.$
@@
This follows from Definitions A.2 and A.18. See [29, Example 5.2.6] for details. ∎
@@
See [30, Corollary 3.1.2] for the proof of (i), [30, Theorem 2.3.3] for the proof of (ii), and [30, Theorem 2.2.1, Theorem 2.2.3, and Theorem 3.2.1.] for the proof of (iii). The proof of (iv) follows by using (i) with $f_{1}=f$ and $f_{2}=g\circ\bm{A}$ and then using (iii). ∎
@@
    $\sup_{\boldsymbol{p}\in\mathbb{R}^{m}}\left\{-f_{1}^{*}(-\bm{A}^{\top}\boldsymbol{p})-f_{2}^{*}(\boldsymbol{p})\right\}$ (48)
@@
See [42, Theorem 31.2 and Corollary 31.2.1] and [3, Proposition 1, Theorem 1 and Theorem 2, pages 163–167] for a proof. The equivalences in (47) and (49) follow from (46) and the rules in Proposition A.2. The equivalence in (50) follows from Fenchel’s inequality (44). ∎
@@
The assumptions in (46) read: there exists $\boldsymbol{x}\in\mathrm{dom~}f_{1}$ such that $\bm{A}\boldsymbol{x}\in\mathrm{ri}\ \mathrm{dom~}f_{2}$ and there exists $\boldsymbol{p}\in\mathrm{dom~}f_{2}^{*}$ such that $\bm{A}^{\top}\boldsymbol{p}\in\mathrm{dom~}f_{1}^{*}$. If the first assumption holds, then Proposition A.2 implies the function $\boldsymbol{x}\mapsto f_{1}(\boldsymbol{x})+f_{2}(\bm{A}\boldsymbol{x})$ is in $\Gamma_{0}(\mathbb{R}^{n})$. Moreover,
@@
Appendix B Technical Proofs
@@
Part 1. First, we show there is $\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})>0$ such that $\left\|{-\bm{A}^{\top}(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}))}\right\|_{\infty}\leqslant 1$ for every $\Delta\in[0,\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})]$. Multiply the vector $\bm{A}^{\top}(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}))$ by the matrix of signs $\bm{D}(\boldsymbol{p}_{0})$ and take the inner product with respect to the unit vector $\boldsymbol{e}_{j}$ with $j\in\{1,\dots,n\}$:
@@
First, suppose $j\in\mathcal{E}(\boldsymbol{p}_{0})$. Then $\langle-\bm{D}(\boldsymbol{p}_{0})\bm{A}^{\top}\boldsymbol{p}_{0},\boldsymbol{e}_{j}\rangle=1$ and $\langle\bm{D}(\boldsymbol{p}_{0})\bm{A}^{\top}\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}),\boldsymbol{e}_{j}\rangle\geqslant 0$ by the KKT condition (17a). If the latter is strictly positive, then we can increase $\Delta$ until the left-hand-side of (51) is equal to 1-1. The smallest such number is
@@
which may be the extended value $\{+\infty\}$ if $\langle\bm{D}(\boldsymbol{p}_{0})\bm{A}^{\top}\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}),\boldsymbol{e}_{j}\rangle=0$ for every $j\in\mathcal{E}(\boldsymbol{p}_{0})$.
@@
is the smallest number for which $\langle\bm{D}(\boldsymbol{p}_{0})\bm{A}^{\top}\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}),\boldsymbol{e}_{j}\rangle\geqslant 0$ among $j\in\mathcal{E}^{\mathsf{C}}(\boldsymbol{p}_{0})$, which may be the extended value $\{+\infty\}$ if $\langle\bm{D}(\boldsymbol{p}_{0})\bm{A}^{\top}\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}),\boldsymbol{e}_{j}\rangle=0$ for every $j\in\mathcal{E}^{\mathsf{C}}(\boldsymbol{p}_{0})$.
@@
Part 3. Let $\Delta\in[0,\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})]$ and $\boldsymbol{q}\in\mathbb{R}^{m}$. Since $\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})\in\partial_{\boldsymbol{p}}V(\boldsymbol{p}_{0};t,\boldsymbol{b})$ and $V(\cdot;t,\boldsymbol{b})$ is $t$-strongly convex, the subdifferentiability property implies
@@
$(1-t\Delta)\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})\in\partial_{\boldsymbol{p}}V(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b});t,\boldsymbol{b})\ \text{for every $\Delta\in[0,\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})]$}.$
@@
$\mathrm{proj}_{\partial_{\boldsymbol{p}}V(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b});t,\boldsymbol{b})}(\boldsymbol{0})\equiv\boldsymbol{d}(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b});t,\boldsymbol{b})=(1-t\Delta)\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}).$
@@
Use Proposition 3.1 with hyperparameter $t_{0}+\delta_{0}$ to get
$\boldsymbol{d}(\boldsymbol{p}_{0};t_{0}+\delta_{0},\boldsymbol{b})=\bm{A}\bm{D}(\boldsymbol{p}_{0})\hat{\boldsymbol{u}}(\boldsymbol{p}_{0};t_{0}+\delta_{0},\boldsymbol{b})-\boldsymbol{b}-(t_{0}+\delta_{0})\boldsymbol{p}_{0},$
@@
$\hat{\boldsymbol{u}}(\boldsymbol{p}_{0};t_{0}+\delta_{0},\boldsymbol{b})\in\operatorname*{arg\,min}_{\boldsymbol{u}\in\mathbb{R}^{n}}\left\|{\bm{A}\bm{D}(\boldsymbol{p}_{0})\boldsymbol{u}-\boldsymbol{b}-(t_{0}+\delta_{0})\boldsymbol{p}_{0}}\right\|_{2}^{2}$
[…]
Now let $\boldsymbol{v}\in\mathbb{R}^{m}$ and use the change of variables $\boldsymbol{u}=\hat{\boldsymbol{u}}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b})+\boldsymbol{v}$. The constraints of the problem become $\boldsymbol{v}_{\mathcal{E}(\boldsymbol{p}_{0})}\geqslant-\hat{\boldsymbol{u}}_{\mathcal{E}(\boldsymbol{p}_{0})}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b})$ and $\boldsymbol{v}_{\mathcal{E}^{\mathsf{C}}(\boldsymbol{p}_{0})}=\boldsymbol{0}$, while the objective function becomes
$\left\|{\bm{A}\bm{D}(\boldsymbol{p}_{0})\boldsymbol{u}-\boldsymbol{b}-(t_{0}+\delta_{0})\boldsymbol{p}_{0}}\right\|_{2}^{2}$
$\qquad\qquad\qquad=\left\|{\bm{A}\bm{D}(\boldsymbol{p}_{0})\hat{\boldsymbol{u}}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b})+\bm{A}\bm{D}(\boldsymbol{p}_{0})\boldsymbol{v}-\boldsymbol{b}-(t_{0}+\delta_{0})\boldsymbol{p}_{0}}\right\|_{2}^{2}$
$\qquad\qquad\qquad=\left\|{\left(\bm{A}\bm{D}(\boldsymbol{p}_{0})\hat{\boldsymbol{u}}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b})-\boldsymbol{b}-t\boldsymbol{p}_{0}\right)+\bm{A}\bm{D}(\boldsymbol{p}_{0})\boldsymbol{v}-\delta_{0}\boldsymbol{p}_{0}}\right\|_{2}^{2}$
$\qquad\qquad\qquad=\left\|{\boldsymbol{d}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b})+\bm{A}\bm{D}(\boldsymbol{p}_{0})\boldsymbol{v}-\delta_{0}\boldsymbol{p}_{0}}\right\|_{2}^{2},$
From this, we obtain (32) and (33). Next, set $\boldsymbol{d}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b})=\boldsymbol{0}$, $\delta_{0}=t-t_{0}$ with $t\in[0,t_{0}]$ and factor out the term $(1-t/t_{0})$ outside the optimization problem to obtain (34) and (35).
B.3 Proof of Proposition 5.1
First, we invoke Lemma 5.1(ii) with
@@
$\boldsymbol{d}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b});t,\boldsymbol{b})=(1-t/t_{0})(\bm{A}\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\hat{\boldsymbol{v}}(t_{0},t,\boldsymbol{b})+t_{0}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))),$ (56)
@@
$\hat{\boldsymbol{v}}(t_{0},t,\boldsymbol{b})\in\operatorname*{arg\,min}_{\boldsymbol{v}\in\mathbb{R}^{n}}\left\|{\bm{A}\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\boldsymbol{v}+t_{0}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}\right\|_{2}^{2}$ (57)
[…]
At $t=t_{0}$, problem (57) reduces to (36) with the identification $\hat{\boldsymbol{v}}(t_{0},t_{0},\boldsymbol{b})\equiv\hat{\boldsymbol{v}}^{s}(t_{0},\boldsymbol{b})$.
We will now identify how the solution $\hat{\boldsymbol{v}}^{s}(t_{0},\boldsymbol{b})$ behaves as we decrease $t_{0}$. There are two potential sources of changes: the set of constraints in (57) and the minimal selection principle via the evolution rule (20) in Proposition 4.1. We turn to these two sources in turn.